3.5 moles of CO2 in a container were expanded against an external pressure of 2.5 atm to a final volume of 37L. The initially pressure and temperature were 8 atm and 30⁰C, respectively. Calculate the final temperature and the work done for the process.


1
Expert's answer
2022-01-25T07:11:22-0500

Gives

"P_2=2.5atm \\n=3.5mol\\V_2=37L"

Ideal gas equation

PV=nRT

"T_2=\frac{P_2V_2}{nR}"

"T_2=\frac{2.5\times37}{3.5\times0.0821}=321.90K"

"T_2=321.90-273=48.9\degree C"

Work

"W=nC_v\u0394T"

"W=3.5\times25.08(321.90-300)"

"W=1922.382J"

W=1.922KJ